# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ 1. Introduction
It is well-known that there are only finitely many elliptic curves over a number field $F$ contained in any $F$-rational isogeny class, and each elliptic curve $E/F$ has a finite torsion subgroup by the Mordell-Weil Theorem. Prior work has investigated how the torsion subgroup varies within an $F$-rational isogeny class; see, for example, [28, 18, 32, 10]. In this paper, we instead study torsion points of elliptic curves within a fixed geometric isogeny class. Here, two elliptic curves lie in the same geometric isogeny class if they are connected by an isogeny defined over $\overline{\mathbb{Q}}$. Any geometric isogeny class contains infinitely many elliptic curves, up to $\overline{\mathbb{Q}}$-isomorphism, and there is no upper bound on the size of a torsion subgroup. Instead, we introduce the problem of studying low degree points associated to $\mathcal{E}$ on modular curves. Motivated by ties to Serre’s Uniformity Problem [8, Theorem 1.2], we focus on the modular curve $X_{1}(N)$, whose non-cuspidal points parametrize elliptic curves with a distinguished point of order $N$. We recall a more precise definition of this curve and the degree of its points in §$\S 2.4$.
For a given geometric isogeny class $\mathcal{E}$ of elliptic curves, our central questions are as follows:
  1. (1)
    For $N\in\mathbb{Z}^{+}$, what is the least degree of a point on $X_{1}(N)$ associated to $\mathcal{E}$?
  2. (2)
    What elliptic curve(s) in $\mathcal{E}$ attain a point on $X_{1}(N)$ of least possible degree?
We call any $E\in\mathcal{E}$ giving a point on $X_{1}(N)$ in least possible degree among those associated to $\mathcal{E}$ a minimal torsion curve for $\mathcal{E}$ of level $N$. To answer these questions it is necessary to determine whether a point of least degree corresponds to an elliptic curve $E$ with $[\mathbb{Q}(j(E)):\mathbb{Q}]$ minimal for $\mathcal{E}$, or whether one must work with an elliptic curve whose $j$-invariant defines an extension of larger degree, but with exceptional arithmetic. In fact both can occur.
@@
Let $\mathcal{E}$ be the $\overline{\mathbb{Q}}$-isogeny class containing an elliptic curve $E/\mathbb{Q}$ with $j(E)=1875$. There is a degree 12 point on $x\in X_{1}(5)$ with $j(x)=j(E)$, and an appropriate twist of $E$ is a minimal torsion curve for $\mathcal{E}$ of level 5. However, no $E^{\prime}\in\mathcal{E}$ with $j(E^{\prime})\in\mathbb{Q}$ is a minimal torsion curve of level $5^{k}$ if $k>1$. The least degree of a point on $X_{1}(5^{k})$ associated to $\mathcal{E}$ is $12\cdot 5^{2k-3}$, and this is attained by $E^{\prime}\in\mathcal{E}$ with $[\mathbb{Q}(j(E^{\prime})):\mathbb{Q}]=6$. More generally, we must move away from rational $j$-invariants to obtain a minimal torsion curve in all known cases where $\mathcal{E}$ contains $E/\mathbb{Q}$ with mod $\ell$ Galois representation having image contained in the normalizer of a Cartan subgroup. This is of interest in the context of Serre’s Uniformity Problem [34]. See Proposition 8.1 for details.
Our first results address the case where $\mathcal{E}$ is rational, i.e., contains an elliptic curve with $j$-invariant in $\mathbb{Q}$. It is natural to expect distinct phenomena if $\mathcal{E}$ consists of elliptic curves with complex multiplication (CM), so we treat non-CM and CM classes separately. For example, the following theorem gives answers to questions (1) and (2) above in the non-CM case upon restriction to points of odd degree, strengthening [8, Proposition 4.1].
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of elliptic curves which is non-CM. If $E\in\mathcal{E}$ and $x=[E,P]\in X_{1}(\ell^{k})$ is a point of odd degree, then $\ell\in\{2,3,5,7,11,13\}$. The least odd degree point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$ is given in Propositions 5.1, 6.1 and 7.1, while the following divisibility conditions hold and are best-possible without placing restrictions on $\mathcal{E}$:
@@
Moreover, among odd degree points on $X_{1}(\ell^{k})$ coming from $\mathcal{E}$, a point of least odd degree can always be associated to $E_{min}\in\mathcal{E}$ with $j(E_{min})\in\mathbb{Q}$ or which is $\ell$-isogenous to an elliptic curve having rational $j$-invariant.
@@
The exceptional class for $\ell=7$ has also been identified in another context. By Sutherland [38], elliptic curves $E/\mathbb{Q}$ with $j(E)=3^{3}\cdot 5\cdot 7^{5}/2^{7}$ provide the only counterexamples over $\mathbb{Q}$ to a local-global principle for rational isogenies of prime degree.
In particular, Theorem 1.1 states that under the given assumptions, there exists a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ which is at most $\ell$-isogenous to an elliptic curve having rational $j$-invariant. Away from points of odd degree, Proposition 8.1 shows the same condition holds for $\mathcal{E}$ containing a rational elliptic curve with $\ell$-adic Galois representation of level $\ell$. These are special cases of the following result, which is a consequence of Serre’s Open Image Theorem [34]. Our proof does not make the constant $C$ explicit, though we suspect such a construction exists; see Remark 4.4.
Proposition 1.3.
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, and let $\ell$ be prime. There exists a constant $C=C(\mathcal{E},\ell)$ such for any $k\in\mathbb{Z}^{+}$ a point of least degree on $X_{1}(\ell^{k})$ coming from $\mathcal{E}$ can be associated to $j_{\text{min}}\in\mathcal{E}$ which is $d$-isogenous to a rational $j$-invariant for some $d\leq C$.
If $\mathcal{E}$ is a class of elliptic curves with complex multiplication, then Serre’s Open Image Theorem does not apply. In this case, $\mathcal{E}$ consists of elliptic curves with CM by various orders within a fixed imaginary quadratic field $K$. Any elliptic curve $E$ with CM by the full ring of integers in $K$ will have $[\mathbb{Q}(j(E)):\mathbb{Q}]$ minimal for $\mathcal{E}$, and we investigate the isogeny distance from $E$ to a minimal torsion curve with $j$-invariant $j_{min}$.
@@
Let $\mathcal{E}$ be a $\overline{\mathbb{Q}}$-isogeny class of elliptic curves with CM by an order in the imaginary quadratic field $K$. For any prime $\ell$, the the least degree of a point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$ is given in Propositions 9.1, 9.2, and 9.3. If $\ell$ is split in $K$, then an elliptic curve with CM by the full ring of integers in $K$ is a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ and $[\mathbb{Q}(j_{min}):\mathbb{Q}]=h_{K}$, the class number of $K$. Otherwise $[\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $k\rightarrow\infty$.
@@
$[\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E/F,\ell^{\infty}}]=[\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0}/F,\ell^{\infty}}].$
See, for example, [25, Proposition 2.1.1]. This can be leveraged to give lower bounds on the degree of a point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$ in terms of $[\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},\ell^{\infty}}]$. We obtain the following proposition, which strengthens [8, Lemma 4.6].
@@
In some cases, we show these lower bounds are best-possible by explicitly constructing $E^{\prime}\in\mathcal{E}$ giving a degree $\delta$ point on $X_{1}(\ell^{k})$. A natural example is when $d=0$ and $\ell$ is odd, in which case a twist of $E_{0}$ is a minimal torsion curve for $X_{1}(\ell^{k})$; if $d>0$, it may be that $j(E^{\prime})\notin\mathbb{Q}$. In other instances, the lower bounds of Proposition 1.5 can be strengthened by showing that a point on $X_{1}(\ell^{k})$ in degree $\delta$ would produce a subgroup of $\operatorname{im}\rho_{E_{0}/\mathbb{Q},\ell^{\infty}}$ which does not occur; see Proposition 6.6. Throughout we make use of the partial classification of images of $\ell$-adic Galois representations of elliptic curves over $\mathbb{Q}$ due to Rouse and Zureick-Brown [33] and Rouse, Sutherland, and Zureick-Brown [27].
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. By Proposition 4.3, any minimal torsion curve for $X_{1}(\ell^{k})$ is at most $C$-isogenous to an elliptic curve with rational $j$-invariant for some $C$ which does not depend on $k$. Thus there exists a finite set of $j$-invariants $\mathcal{J}=\{j_{1},j_{2},\dots,j_{r}\}$ such that for any $k\in\mathbb{Z}^{+}$, there exists a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ with $j$-invariant in $\mathcal{J}$. However, our proof does not make $\mathcal{J}$ explicit, so our results may not be obtained by checking a finite number of $j$-invariants. On the other hand, for a fixed $k\in\mathbb{Z}^{+}$, a finite check is possible since the $j$-invariant of any minimal torsion curve must generate an extension degree at most $\deg(X_{1}(\ell^{k})\rightarrow X_{1}(1))$.
Results concerning CM isogeny classes rely heavily on prior work of the first author in collaboration with Clark [3, 4].
@@
Cremona and Najman [14] prove numerous results concerning torsion points of $\mathbb{Q}$-curves defined over number fields of odd degree. Any such elliptic curve is necessarily isogenous to one having rational $j$-invariant, providing immediate ties to our study of minimal torsion curves in rational geometric isogeny classes. This class of $\mathbb{Q}$-curves is again studied in [8], where the first author and Najman show the $\overline{\mathbb{Q}}$-isogeny class containing the elliptic curve with $j$-invariant $-140625/8$ is the unique rational non-CM class giving rise to a sporadic point of odd degree on any modular curve $X_{1}(N)$. More recent work of Genao [21] provides “typical” bounds on the size of the torsion subgroup of an elliptic curve over a number field which belongs to a rational $\overline{\mathbb{Q}}$-isogeny class, and his subsequent work [20] pursues polynomial bounds on such torsion subgroups.
Prior work of the first author and Clark [4] gives the least degree of a point on $X_{1}(N)$ associated to an elliptic curve with CM by a fixed order in an imaginary quadratic field, while work of Clark, Genao, Pollack, and Saia [12] investigates the least degree of any CM point on $X_{1}(N)$. Our results fall somewhat in the middle of these two directions of study – we investigate the least degree of a CM point across all orders within a fixed imaginary quadratic field.

### v2

@@ 1. Introduction
The modular curve $X_{1}(N)$ is an algebraic curve over $\mathbb{Q}$ whose non-cuspidal points parameterize elliptic curves with a distinguished point of order $N$. Understanding degree $d$ points [^1] Recall that the degree of a closed point $x$ on a curve $C$ over a number field $k$ is the degree of its residue field: $[k(x):k]$. This is equivalent to the size of the Galois orbit of points in $C(\overline{k})$ associated to $x$. on $X_{1}(N)$ is an essential step in classifying torsion subgroups of elliptic curves defined over number fields of degree $d$ — a problem which is solved only for $d\leq 4$ [mazur77, kamienny86, KM88, kamienny92, Deg3Class, classificationdeg4points]. By the Riemann-Roch Theorem, the curve $X_{1}(N)$ has infinitely many points in any degree greater than its genus, so it suffices to characterize points of low degree. One important class of low degree points are sporadic points, which are $x\in X_{1}(N)$ for which there are only finitely many points of degree at most $\deg(x)$. Ruling out unexpected sporadic points on $X_{1}(N)$ would have implications for major open questions in the field, including Serre’s Uniformity problem (as in [OddDegQCurve, Theorem 1.3]) and uniformity conjectures for the modular curve $X_{0}(N)$; see [ogg, Conjecture 18] and [QuadPointsConjecture2024, Conjecture 1.1]. Such applications further motivate the need for an improved understanding of low degree points
While studying low degree points on $X_{1}(N)$ is quite difficult in general, the problem can be made more tractable by restricting the class of points under consideration. One line of investigation considers only points associated to elliptic curves with complex multiplication (CM); see, for example, [BC2, LeastCMdegree, ClarkVolcanoes]. The least degree of a point on $X_{1}(N)$ associated to an elliptic curve with CM by an order $\mathcal{O}$ in an imaginary quadratic field $K$ is given in [BC2, Theorem 1.2]. One could more generally hope to characterize the least degree of a point on $X_{1}(N)$ as we range over elliptic curves with CM by any order in $K$. Since an elliptic curve with CM by $\mathcal{O}\subseteq K$ is isogenous over $\overline{\mathbb{Q}}$ to one with CM by the full ring of integers, this is in essence a special case of the following question: For a given geometric isogeny class $\mathcal{E}$ of elliptic curves and fixed $N\in\mathbb{Z}^{+}$, what is the least degree of a point on $X_{1}(N)$ associated to an elliptic curve in $\mathcal{E}$? And what elliptic curve(s) in $\mathcal{E}$ attain a point on $X_{1}(N)$ of least possible degree? We investigate these questions in the present work.
Let $\mathcal{E}$ be a geometric isogeny class of elliptic curves. That is, for any $E_{1},E_{2}\in\mathcal{E}$, there exists an isogeny $\varphi:E_{1}\rightarrow E_{2}$ defined over $\overline{\mathbb{Q}}$. We want to find the minimum $d$ such that there exists $F$ of degree $d$ and $E/F\in\mathcal{E}$ such that $E(F)$ contains a point $P$ of order $N$. We call $E$ a minimal torsion curve for $\mathcal{E}$ of level $N$. Because $F$ is of minimal degree, the residue field of the closed point associated to $(E,P)\in X_{1}(N)(\overline{\mathbb{Q}})$ is also of degree $d$, and so we may alternatively view $E$ as a curve over $\overline{\mathbb{Q}}$ identified by its $j$-invariant, $j_{min}$. In this paper, we study minimal torsion curves of prime-power level. For CM isogeny classes, we obtain a near complete characterization.
Theorem 1.1.
Let $\mathcal{E}$ be a $\overline{\mathbb{Q}}$-isogeny class of elliptic curves with CM by an order in the imaginary quadratic field $K$. For a prime number $\ell$, the the least degree of a point on $X_{1}(\ell^{k})$ associated to any $E\in\mathcal{E}$ is given in Propositions 9.1, 9.2, and 9.3. If $\ell$ is split in $K$, then an elliptic curve with CM by the full ring of integers in $K$ is a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ and $\mathbb{Q}(j_{min}):\mathbb{Q}]=h_{K}$, the class number of $K$. Otherwise $\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $k\rightarrow\infty$.
If $\mathcal{E}$ is a $\overline{\mathbb{Q}}$-isogeny class of elliptic curves with CM by an order in $K$, then $\mathbb{Q}(j(E)):\mathbb{Q}]\geq h_{K}$ for any $E\in\mathcal{E}$. Thus when $\ell$ is split in $K$, a minimal torsion curve has $\mathbb{Q}(j_{min}):\mathbb{Q}]$ minimal for $\mathcal{E}$. However, if $\ell$ is inert or ramified in $K$, the extension $\mathbb{Q}(j_{min}):\mathbb{Q}]$ will generally not be minimal for $\mathcal{E}$, and indeed must grow with $k$. These minimal torsion curves have exceptional arithmetic, such as the existence of a high-power $\ell^{k}$-isogeny defined over $\mathbb{Q}(j_{min})$, which more than compensates for the degree of $\mathbb{Q}(j_{min})$. This phenomenon was previously observed for CM points of odd degree in [BP, Remark 2.7]. We illustrate with the following example.
@@
Let $\mathcal{E}$ be the $\overline{\mathbb{Q}}$-isogeny class of $E/\mathbb{Q}$ with LMFDB label 27.a1. Here $K=\mathbb{Q}(\sqrt{-3})$, and we take $\ell=3$ which is ramified in $K$. There are two rational $j$-invariants associated to elliptic curves in $\mathcal{E}$, namely 12288000-12288000 and 0. Thus $1$ is the minimal degree of an extension generated by the $j$-invariant of $E\in\mathcal{E}$. However, by Proposition 9.3, the $j$-invariant of a minimal torsion curve of level $3^{k}$ generates an extension of degree at least $3^{(k-5)/2}$. Thus there is no minimal torsion curve with rational $j$-invariant for $k\geq 6$, and the degree of $\mathbb{Q}(j_{min})$ must grow with $k$.
Even in non-CM isogeny classes, minimal torsion curves may have $j$-invariant generating a field of unexpectedly large degree. In some cases, these minimal torsion curves improve degree bounds arising from elliptic curves having rational $j$-invariant.
Example 2.
Let $\mathcal{E}$ be the $\overline{\mathbb{Q}}$-isogeny class containing $E/\mathbb{Q}$ with LMFDB label 9225.l1. Here $j(E)\in\mathbb{Q}$, but in Section 88 we see any minimal torsion curve for $\mathcal{E}$ of level 49 has $\mathbb{Q}(j_{min}):\mathbb{Q}]\geq 2$. This class is notable since $\mathcal{E}$ contains an elliptic curve giving a point on $X_{1}(49)$ of degree $42$, which is lower than the degree of any $x\in X_{1}(49)$ with $j(x)\in\mathbb{Q}$. See Corollary 8.3.
Our next results concern the case where $\mathcal{E}$ is rational, i.e., contains an elliptic curve with $j$-invariant in $\mathbb{Q}$. For non-CM classes, we obtain a characterization of minimal torsion curves upon restriction to points of odd degree, strengthening [OddDegQCurve, Proposition 4.1].
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of elliptic curves which is non-CM. If $E\in\mathcal{E}$ and $x=[E,P]\in X_{1}(\ell^{k})$ is a point of odd degree, then $\ell\in\{2,3,5,7,11,13\}$. The least odd degree point on $X_{1}(\ell^{k})$ associated to $E^{\prime}\in\mathcal{E}$ is given in Propositions 5.1, 6.1 and 7.1, while the following divisibility conditions hold and are best-possible across all such $\mathcal{E}$:
@@
Moreover, among odd degree points on $X_{1}(\ell^{k})$ coming from $E^{\prime}\in\mathcal{E}$, a point of least odd degree can always be associated to $E_{min}\in\mathcal{E}$ with $j(E_{min})\in\mathbb{Q}$ or which is $\ell$-isogenous to an elliptic curve having rational $j$-invariant.
@@
The exceptional class for $\ell=7$ has also been identified in another context. By Sutherland [Sutherland12], elliptic curves $E/\mathbb{Q}$ with $j(E)=3^{3}\cdot 5\cdot 7^{5}/2^{7}$ provide the only counterexamples over $\mathbb{Q}$ to a local-global principle for rational isogenies of prime degree.
In particular, Theorem 1.2 states that under the given assumptions, there exists a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ which is at most $\ell$-isogenous to an elliptic curve having rational $j$-invariant. Away from points of odd degree, Proposition 8.1 shows the same condition holds for $\mathcal{E}$ containing a rational elliptic curve with $\ell$-adic Galois representation of level $\ell$. These are special cases of the following result, which is a consequence of Serre’s Open Image Theorem [serre72].
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, and let $\ell$ be prime. There exists a constant $C=C(\mathcal{E},\ell)$ such that for any $k\in\mathbb{Z}^{+}$ a point of least degree on $X_{1}(\ell^{k})$ coming from $E^{\prime}\in\mathcal{E}$ can be associated to $j_{min}\in\mathcal{E}$ which is $d$-isogenous to a rational $j$-invariant for some $d\leq C$.
Thus unlike the CM case addressed in Theorem 1.1, there always exists a minimal torsion curve of level $\ell^{k}$ with $j$-invariant in an extension of bounded degree. Unfortunately, our proof does not make the constant $C$ explicit; see Theorem 4.3. If $\mathcal{E}$ contains a non-CM elliptic curve $E/\mathbb{Q}$ with $\ell$-adic Galois representation of level $\ell$, then $C=\ell$ by Proposition 8.1. However, we fail to identify a more general connection between $C$ and the level of $E/\mathbb{Q}\in\mathcal{E}$, though it is natural to ask whether one exists (Remark 4.4).
@@
$\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E/F,\ell^{\infty}}]=[\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0}/F,\ell^{\infty}}].$
See, for example, [greenberg2012, Proposition 2.1.1]. This can be leveraged to give lower bounds on the degree of a point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$ in terms of $\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},\ell^{\infty}}]$. We obtain the following proposition, which strengthens [OddDegQCurve, Lemma 4.6].
@@
In some cases, we show these lower bounds are best-possible by explicitly constructing $E^{\prime}\in\mathcal{E}$ giving a degree $\delta$ point on $X_{1}(\ell^{k})$. A natural example is when $d=0$ and $\ell$ is odd, in which case a twist of $E_{0}$ is a minimal torsion curve for $X_{1}(\ell^{k})$; if $d>0$, it may be that $j(E^{\prime})\notin\mathbb{Q}$. In other instances, the lower bounds of Proposition 1.5 can be strengthened by showing that a point on $X_{1}(\ell^{k})$ in degree $\delta$ would produce a subgroup of $\operatorname{im}\rho_{E_{0}/\mathbb{Q},\ell^{\infty}}$ which does not occur; see Proposition 6.6. Throughout we make use of the partial classification of images of $\ell$-adic Galois representations of elliptic curves over $\mathbb{Q}$ due to Rouse and Zureick-Brown [RouseDZB] and Rouse, Sutherland, and Zureick-Brown [RSZ21].
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. By Theorem 4.3, any minimal torsion curve for $X_{1}(\ell^{k})$ is at most $C$-isogenous to an elliptic curve with rational $j$-invariant for some $C$ which does not depend on $k$. Thus there exists a finite set of $j$-invariants $\mathcal{J}=\{j_{1},j_{2},\dots,j_{r}\}$ such that for any $k\in\mathbb{Z}^{+}$, there exists a minimal torsion curve for $\mathcal{E}$ of level $\ell^{k}$ with $j$-invariant in $\mathcal{J}$. However, our proof does not make $\mathcal{J}$ explicit, so our results may not be obtained by checking a finite number of $j$-invariants. On the other hand, for a fixed $k\in\mathbb{Z}^{+}$, a finite check is possible since the $j$-invariant of any minimal torsion curve must generate an extension degree at most $\deg(X_{1}(\ell^{k})\rightarrow X_{1}(1))$.
Results concerning CM isogeny classes rely heavily on prior work of the first author in collaboration with Clark [BC1, BC2].
@@
Cremona and Najman [CremonaNajmanQCurve] prove numerous results concerning torsion points of $\mathbb{Q}$-curves defined over number fields of odd degree. Any such elliptic curve is necessarily isogenous to one having rational $j$-invariant, providing immediate ties to our study of minimal torsion curves in rational geometric isogeny classes. This class of $\mathbb{Q}$-curves is again studied in [OddDegQCurve], where the first author and Najman show the $\overline{\mathbb{Q}}$-isogeny class containing the elliptic curve with $j$-invariant 140625/8-140625/8 is the unique rational non-CM class giving rise to a sporadic point of odd degree on any modular curve $X_{1}(N)$. More recent work of Genao [Genao21] provides “typical” bounds on the size of the torsion subgroup of an elliptic curve over a number field which belongs to a rational $\overline{\mathbb{Q}}$-isogeny class, and his subsequent work [Genao22] pursues polynomial bounds on such torsion subgroups.
Prior work of the first author and Clark [BC2] gives the least degree of a point on $X_{1}(N)$ associated to an elliptic curve with CM by a fixed order in an imaginary quadratic field, while work of Clark, Genao, Pollack, and Saia [LeastCMdegree] investigates the least degree of any CM point on $X_{1}(N)$. Our results fall somewhat in the middle of these two directions of study – we investigate the least degree of a CM point across all orders within a fixed imaginary quadratic field.
1.3. Code.
We make frequent use of the computer algebra system Magma [Magma]. All code is available at https://github.com/abbey-bourdon/minimal_torsion_curves.